Hi i am really hoping you can help me here, i have never uses a tutor before but i was frantically trying to find something to help me so here it goes...
if i have 3 ordered pairs (3,15) (12,13) (5,7) on a graph and are making a scalene triangle. The question is how to do make this into in the an equation of two of the perpendicular Bisectors. and then getting that information i am not sure how to solve it. i am thinking if i get it started i will be able to finish it but i have NO idea even how to start to get an equation set up.
THANK YOU SO MUCH!

Answer by stanbon(57387)

i have 3 ordered pairs (3,15) (12,13) (5,7) on a graph and are making a scalene triangle. The question is how to do make this into in the an equation of two of the perpendicular Bisectors.
---------------
Choose two of the points: (3,15),(12,13)
Find the slope and midpoint of the line segment joining the two points.
slope: [13-15]/[12-3]= -2/9
mid-pt: x=(3+12)/2=15/2
........y=(13+15)/2=14
Now find the equation of the line perdicular to the line segment and
passing thru the mid-p.
The slope of the bisector is 9/2 because it must be the negative inverse
of the line segment;s slope.
---------
So, slope = 9/2 and it passes thru (15/2,14)
Form is y=mx+b
Substitute the x,y, and m values to find "b", as follows:
14=(9/2)(15/2)+b
14=135/4 + b
56/4 = 135/4+b
b=-79/4
EQUATION of one of the perpendicular bisectors: y=(9/2)x-79/4
------------
Follow the same procedure to find the 2nd perpendicular bisector.
============
Cheers,
Stan H.